# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

### v1

@@ Abstract
The rapid development of multimodal large reasoning models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 11 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, a long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to addressing safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhances the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
@@ 1 Introduction
The development of large reasoning models (LRMs) such as OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) has unveiled significant potential in addressing complex reasoning tasks Chan et al. (2024); Chen et al. (2024), including mathematical problem-solving Shao et al. (2024) and code generation El-Kishky et al. (2025). Consequently, the reasoning paradigm has been naturally extended to multimodal scenarios, giving rise to multimodal large reasoning models (MLRMs) Peng et al. (2025a); Yao et al. (2024); Team et al. (2025). The MLRMs are capable of generating long reasoning pathways and exhibit remarkable spatial reasoning abilities, thereby broadening their real-world applicability across diverse scenarios Zhang et al. (2025a); Huang et al. (2025); Wu et al. (2025).
@@
While LRMs have led to a transformative leap in reasoning performance, prior studies have indicated that the exposure of the chain-of-thought (CoT) process inadvertently undermines safety, as adversaries exploit intermediate reasoning steps to circumvent alignment safeguards Jiang et al. (2025); Zhou et al. (2025); Ying et al. (2025). This vulnerability has also been observed in MLRMs Fang et al. (2025). However, current research on the safety of MLRMs remains limited, primarily focusing on two jailbreak robustness benchmarks, with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
As shown in Figure 1, we categorize the safety benchmarks of multimodal models into two key aspects: safety-awareness and jailbreak robustness. Jailbreak robustness benchmarks focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms. Safety-awareness benchmarks emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications. Compared to unimodal LRMs, MLRMs exhibits the following research questions:
@@
RQ2: What are the limitations of incorporating additional modalities in MLRMs for safety concerns?
@@
To investigate the above research questions and challenges, in this study, we first conduct a systematic safety evaluation of advanced MLRMs, including Kimi-VL-thinking Team et al. (2025), R1-Onevision Yang et al. (2025) and etc. The results demonstrate that MLRMs have significant negative impact on the safety performance, while the performance degradation is task-dependent. In particular, MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs. Motivated by this, we further propose a data construction method that incorporates safety-oriented thought process to investigate the effectiveness of this insight.
@@
  • This study conducts a systematic evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue of incorporating additional modalities.
  • Experimental results demonstrate that our method improve the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
Model Name Base Model
LlamaV-o1 Thawakar et al. (2025) Llama3.2-vision-11B
Llava-cot Xu et al. (2024) Llama3.2-vision-11B
Mulberry-Llama Yao et al. (2024) Llama3.2-vision-11B
LMM-R1 Peng et al. (2025b) Qwen2.5-VL-3B
R1-Onevision Yang et al. (2025) Qwen2.5-VL-7B
MM-Eureka Meng et al. (2025) InternVL-2.5-8B
Kimi-VL-thinking Team et al. (2025) Kimi-VL-instruct
Skywork-R1V Peng et al. (2025a) -
QVQ-Preview -
Gemini2.0-flash-thinking -
Claude3.7-sonnet-thinking -
@@
Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 51.09
R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 49.69(-1.40)
@@ 2.1 Evaluation Settings
To comprehensively assess the safety performance of MLRMs across diverse scenarios, we adopt benchmark datasets from two distinct perspectives: safety-awareness and jailbreak robustness. For the assessment of safety-awareness, we employ SIUO Wang et al. (2025) and MSSBench Zhou et al. (2024) datasets, while for evaluating jailbreak robustness, we employ MMSafetyBench Liu et al. (2024b), FigStep Gong et al. (2025) and JailBreaKV Luo et al. (2024).
In safety-awareness tasks, models need to reason jointly over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of malicious crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze the results separately. Further details regarding the datasets are provided in Appendix A.
@@
We evaluate a total of 11 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Table 1. For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments were conducted using four NVIDIA A100-80G GPUs.
Figure 2: Case study of the better safety consideration on safety-awareness tasks. Kimi-VL-Instruct directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities.
@@
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A. Following the settings of MMSafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
@@ 2.2 Safety Results and Inspection
Finding 2 ($\Rightarrow$ RQ1): MLRM paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
@@
As shown in Table 2, the safety score of models such as Kimi-VL-thinking Team et al. (2025), LMM-R1 Peng et al. (2025b), and MM-Eureka Meng et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in attention patterns according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
@@
Qwen2.5-VL-7B (I) 66.49 12.5 29.94
Qwen2.5-VL-7B (T) 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
R1-Onevision (I) 79.57 32.14 17.31
R1-Onevision (T) 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
Kimi-VL-instruct (I) 47.74 22.50 25.00
Kimi-VL-instruct (T) 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
Kimi-VL-thinking (I) 61.49 33.93 35.93
Kimi-VL-thinking (T) 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
Llama3.2-vision-11B (I) 38.45 5.71 37.13
Llama3.2-vision-11B (T) 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
Llava-cot (I) 72.26 33.57 26.95
Llava-cot (T) 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
InternVL2.5-8B (I) 59.64 15.00 28.14
InternVL2.5-8B (T) 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
MM-Eureka (I) 60.12 11.79 28.14
MM-Eureka (T) 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
Table 3: Safety performance when converting image into text caption. “I” indicates multimodal inputs, while “T” refers to text-only inputs.
@@
Figure 4: ASR scores of the thought process and the final response generated by MLRMs on MMSafetyBench respectively.
@@
To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process. As shown in Figure 4, models that explicitly exhibit long CoT Kojima et al. (2022) generate harmful content during the thought process. Additionally, the ASR results for the four models—QVQ-Preview, R1-Onevision, Llava-cot, and Skywork-R1V—are higher for the thought process than for the final responses. Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe responses. As illustrated in Figure 3, although the model detects harmful intent in some cases, it does not reject the prompt outright. Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content. It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
@@
Figure 5: Overview of our data construction. We propose a multi-step pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
@@
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated after reducing the thinking steps.
R1-Onevision Safe Response Unsafe Response
Safe Thought 12.65 1.90
Unsafe Thought 7.34 78.11
Kimi-VL-thinking Safe Response Unsafe Response
Safe Thought 12.68 2.74
Unsafe Thought 25.83 58.75
Llava-cot Safe Response Unsafe Response
Safe Thought 15.89 3.06
Unsafe Thought 11.85 69.23
QVQ-Preview Safe Response Unsafe Response
Safe Thought 24.52 6.91
Unsafe Thought 6.19 62.38
Skywork-R1V Safe Response Unsafe Response
Safe Thought 19.88 3.93
Unsafe Thought 7.44 68.75
@@
Table 4: Proportion of safety in the thought process and the final response on MMSafetyBench.
Comparative Effect between Thought and Response.
@@
As shown in Table 4, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final response. For example, the safe thought process leads to the unsafe response in QVQ-Preview, with respect to 6.91% proportion. In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response. The inconsistency between safe thought and unsafe response is due to the potential security risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final response. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt. In particular, Kimi-VL-thinking demonstrates a relatively strong capacity for self-correction with respect to 25.83% proportion, which aligns with the observations reported in Finding 2. The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.

### v2

@@ Abstract
The rapid development of Multimodal Large Reasoning Models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 11 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, the long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to address safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhances the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
@@ 1 Introduction
The development of Large Reasoning Mode-ls (LRMs) such as OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025) has unveiled significant potential in addressing complex reasoning tasks Chan et al. (2024); Chen et al. (2024), including mathematical problem-solving Shao et al. (2024) and code generation El-Kishky et al. (2025). Consequently, the reasoning paradigm has been naturally extended to multimodal scenarios, giving rise to Multimodal Large Reasoning Models (MLRMs) Peng et al. (2025a); Yao et al. (2024); Team et al. (2025). The MLRMs are capable of generating long reasoning pathways and exhibit remarkable spatial reasoning abilities, thereby broadening their real-world applicability across diverse scenarios Zhang et al. (2025a); Huang et al. (2025); Wu et al. (2025).
@@
While LRMs have led to a transformative leap in reasoning performance, prior studies have indicated that the exposure of the Chain-of-Thought (CoT) process inadvertently undermines safety, as adversaries exploit intermediate reasoning steps to circumvent alignment safeguards Jiang et al. (2025); Zhou et al. (2025); Ying et al. (2025). This vulnerability has also been observed in MLRMs Fang et al. (2025). However, current research on the safety of MLRMs remains limited, primarily focusing on jailbreak robustness benchmarks Gong et al. (2025); Liu et al. (2024b), with insufficient exploration of the broader spectrum of safety challenges in multimodal settings.
As shown in Figure 1, we categorize the safety benchmarks of multimodal models into two key aspects: safety-awareness and jailbreak robustness. Jailbreak robustness benchmarks Gong et al. (2025); Liu et al. (2024b); Luo et al. (2024) focus on evaluating the resilience of the model against deliberately crafted or modified textual prompts and visual inputs that aim to bypass established safety defense mechanisms. Safety-awareness benchmarks Wang et al. (2025); Zhou et al. (2024) emphasize the capability of models to proactively identify potential safety risks embedded in user inputs, aligning more closely with the complex and dynamic safety demands encountered in real-world applications. Compared to unimodal LRMs, MLRMs exhibits the following research questions:
@@
RQ2: What are the risks of incorporating additional modalities in MLRMs for safety concerns?
@@
To investigate the above research questions and challenges, in this study, we first conduct a systematic safety evaluation of advanced MLRMs, including Kimi-VL-Thinking Team et al. (2025), R1-Onevision Yang et al. (2025) , etc. The results demonstrate that MLRMs have significant negative impact on the safety performance, while the performance degradation is task-dependent. Furthermore, we found that MLRMs could identify more potential safety risks through deliberate thinking, leading to higher safety scores on safety-awareness benchmarks, which provides a novel perspective for mitigating safety degradation in MLRMs. Motivated by this, we further propose a data construction method that incorporates safety-oriented thought process to investigate the effectiveness of this insight.
@@
  • This study conducts a systematic safety evaluation of MLRMs and investigates the empirical results, revealing several novel findings and providing new perspectives for the development of safer MLRMs.
  • We construct a multimodal fine-tuning dataset with safety-oriented thought process for safety alignment, alleviating the issue associated with incorporating additional modalities.
  • Experimental results demonstrate that our method improves the safety performance of MLRMs across multiple benchmarks by enabling self-correction thinking along the reasoning pathways, compared with previous defense methods.
@@
Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 50.02
R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 48.94(-1.08)
@@ 2.1 Evaluation Settings
To comprehensively assess the safety performance of MLRMs across diverse scenarios, we adopt benchmark datasets from two distinct perspectives: safety-awareness and jailbreak robustness. For assessment of safety-awareness, we employ SIUO Wang et al. (2025) and MSSBench Zhou et al. (2024) datasets, while for evaluating jailbreak robustness, we employ MM-SafetyBench Liu et al. (2024b), FigStep Gong et al. (2025) and JailBreaKV Luo et al. (2024).
In safety-awareness tasks, models need to jointly reason over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of maliciously crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze these results separately. Further details regarding the datasets are provided in Appendix A.1.
@@
We evaluate a total of 11 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-Thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Appendix A.3. For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments are conducted using four NVIDIA A100-80G GPUs.
@@
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A.1. Following the settings of MM-SafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
Figure 2: Case study of the better safety consideration on safety-awareness tasks. Kimi-VL directly outputs the answer that ignores the potential risk, while Kimi-VL-Thinking dives deeper into the insidious safety issue with stronger reasoning abilities. The red indicates the unsafe parts, while the green indicates the content related to potential risks identified by reasoning models.
@@ 2.2 Safety Results and Inspection
Finding 2 ($\Rightarrow$ RQ1): MLRMs paradoxically enhances safety performance via long reasoning on safety-awareness benchmarks, due to improved capability in identifying unsafe intent.
@@
As shown in Table 1, the safety score of models such as Kimi-VL-Thinking Team et al. (2025), LMM-R1 Peng et al. (2025b), and MM-Eureka Meng et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
@@
Qwen2.5-VL-7B 66.49 12.5 29.94
     -w/o image 48.21(-18.28) 6.79(-5.71) 38.10(+8.16)
R1-Onevision 79.57 32.14 17.31
     -w/o image 65.02(-14.55) 23.21(-8.93) 23.35(+6.04)
Kimi-VL 47.74 22.50 25.00
     -w/o image 57.56(+9.82) 19.29(-3.21) 25.00(+0.00)
Kimi-VL-Thinking 61.49 33.93 35.93
     -w/o image 55.71(-5.78) 29.64(-4.29) 29.34(-6.59)
Llama 3.2-vision-11B 38.45 5.71 37.13
     -w/o image 58.52(+20.07) 44.29(+38.58) 22.50(-14.63)
LLaVA-CoT 72.26 33.57 26.95
     -w/o image 60.06(-12.20) 40.36(+6.79) 28.14(+1.19)
InternVL2.5-8B 59.64 15.00 28.14
     -w/o image 48.57(-11.07) 9.64(-5.36) 27.55(-0.59)
$M$-Eureka 60.12 11.79 28.14
     -w/o image 50.24(-9.88) 10.71(-1.08) 31.14(+3.00)
Table 2: Safety performance when converting image into text caption. The safety performance variation has been marked in brackets, where the red color represents safety deterioration and green stands for safety improvement. MMSafe is the abbreviation for $M$-SafetyBench.
@@
Figure 4: ASR scores of the thought process and the final answer generated by MLRMs on FigStep respectively.
@@
To investigate the impact of reasoning pathways in MLRMs, we independently evaluate the safety performance of the thought process. As shown in Figure 4, the proportion of unsafe thought process generated by the four models (R1-Onevision, LLaVA-CoT, Kimi-VL-Thinking and Skywork-R1V) all exceeds 50 percent. Moreover, we analyze the thought process to identify how intermediate content contributes to the production of unsafe answers. As illustrated in Figure 3, although the model detects harmful intent in some cases, it does not reject the prompt outright. Instead, MLRM engages in self-deception by reframing the intent of users as educational, humorous, or otherwise benign, which rationalizes the model to generate harmful content. It is a critical challenge to address subtle intent rationalization in long reasoning, which needs to be explored in future work.
@@
Figure 5: Overview of our data construction pipeline. We propose a multi-stage pipeline to build the datasets based on various safety-related topics (Step 1: Word Cloud in Blue Block) and image description (Step 2: Red Block), which provides a thoughtful consideration of reasoning with safety, explicitly incorporating long CoT reasoning into the addressing process (Step 3: Green Block) and meticulously designed filtering mechanism (Step 4: Yellow Block).
@@
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is secure.
  R1-Onevision Safe Answer Unsafe Answer
Safe Thought 13.60% 29.40%
Unsafe Thought 14.20% 42.80%
Kimi-VL-Thinking Safe Answer Unsafe Answer
Safe Thought 8.40% 30.40%
Unsafe Thought 4.60% 48.20%
LLaVA-CoT Safe Answer Unsafe Answer
Safe Thought 8.80% 30.20%
Unsafe Thought 6.40% 54.60%
QVQ-Preview Safe Answer Unsafe Answer
Safe Thought 24.60% 46.60%
Unsafe Thought 4.40% 24.20%
Skywork-R1V Safe Answer Unsafe Answer
Safe Thought 9.40% 37.60%
Unsafe Thought 4.80% 48.20%
@@
Table 3: Proportion of safety in the thought process and the final answer on FigStep.
Comparative Effect between Thought and Answer.
@@
As shown in Table 3, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final answer. For example, the safe thought process leads to the unsafe answer in QVQ-Preview, with respect to 46.60% proportion. In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response. The inconsistency between safe thought and unsafe answer is due to the potential security risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final answer. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt. The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.